OpenStudy (marcelie):

find the area enclosed by the x axis and the curve x= t^3+1 y= 2t-t^2

OpenStudy (marcelie):

@zepdrix

OpenStudy (marcelie):

how do i find my bounds ? i totally forgot lol

zepdrix (zepdrix):

For parametric equations? Hmm that's a good question :d

OpenStudy (marcelie):

yeah :O

zepdrix (zepdrix):

One of the boundaries is the x-axis... The x-axis is the line y=0. So our curve will intersect the x-axis when y=0.\[\large\rm y=2t-t^2\]\[\large\rm 0=2t-t^2\]Solve for t.

OpenStudy (marcelie):

i got t= 0 and t= -2

zepdrix (zepdrix):

-2? Hmm that doesn't seem right :d

zepdrix (zepdrix):

Check again lady. t(2-t)=0

OpenStudy (marcelie):

oh wait lmao t= 0 and t= 2

zepdrix (zepdrix):

Okieeee, so those are probably the boundaries we'll use. I'm trying to remember how to find area when given a parameter like this... sec hold upppp :)

OpenStudy (marcelie):

okay to find our bounds we solve for t?

zepdrix (zepdrix):

Yes. That's what we did.

OpenStudy (marcelie):

okie lol

zepdrix (zepdrix):

So I guess our formula for area is,\[\large\rm A=\int\limits_{t_1}^{t_2}y~dx\]We found our boundary values,\[\large\rm A=\int\limits\limits_{0}^{2}y~dx\]And we know what y is,\[\large\rm A=\int\limits\limits\limits_{0}^{2}(2t-t^2)~dx\]Let's try to find the differential of x.

zepdrix (zepdrix):

\[\large\rm x=t^3+1\]\[\large\rm dx=?\]

OpenStudy (marcelie):

3t^2

zepdrix (zepdrix):

Yes,\[\large\rm dx=3t^2~dt\]So therefore our integral is,\[\large\rm A=\int\limits\limits\limits\limits_{0}^{2}(2t-t^2)(3t^2)dt\]

zepdrix (zepdrix):

*I think*

OpenStudy (marcelie):

oh okay so then we take one derivative right ?

zepdrix (zepdrix):

We already did. \(\large\rm x=t^3+1\qquad\to\qquad dx=3t^2~dt\)

OpenStudy (marcelie):

oh yeahh lool

OpenStudy (marcelie):

i got 24/5

zepdrix (zepdrix):

Mmmm ya that's what Wolfram says also :O

zepdrix (zepdrix):

So hopefully we did that properly :d

zepdrix (zepdrix):

ʕっ•ᴥ•ʔっ Bear says, "good job!"

OpenStudy (marcelie):

loool

zepdrix (zepdrix):

ʕ´• ᴥ •`ʔ